A manufacturing company receives orders for engines from two assembly plants. Plant I needs at least 45 ​engines, and plant II needs at least 32 engines. The company can send at most 120 engines to these assembly plants. It costs ​$30 per engine to ship to plant I and ​$40 per engine to ship to plant II. Plant I gives the manufacturing company ​$20 in rebates toward its products for each engine they​ buy, while plant II gives similar ​$10 rebates. The manufacturer estimates that they need at least ​$1500 in rebates to cover products they plan to buy from the two plants. How many engines should be shipped to each plant to minimize shipping​ costs? What is the minimum​ cost?

Answer by Theo(13342)

your objective function is to minimize the cost.

set x equal to the number of engines shipped to plant I.
set y equal to the number of engines shipped to plant II.

total cost = 30x + 40y

that's the function you want to minimize.
it costs 30 dollars to ship 1 engine to plant I.
it costs 40 dollars to ship 1 engine to plant II.

set c = total cost.

your objective function is c = 30x + 40y.

your constraints are:

x >= 45
y >= 32

this means you need to ship at least 45 engines to plant I and at least 32 engines to plant II.

x >= 0
y >= 0

this means that the number of engines shipped has to be equal to or greater than 0.

20x + 10y >= 1500

this means that the total rebates provided by the assembly plants has to be greater than or equal to 1500.

x + y < = 120

this means that the total number of engines that can be shipped has to be less than or equal to 120.

summarize your formulas:

objective function:

c = 30x + 40y

constraints:

x + y <= 120
x >= 45
y >= 32
20x + 10y >= 1500


graph the constraint equations and then shade the areas on the graph that satisfy the inequalities.

this is what my graph looks like.

$$$

i graphed the equations.
i then shaded the areas that satisfied the inequalities.

then you need to find the corner points of the shaded area.

those are:

(45,75)
(45,60)
(59,32)
(88,32)

you then analyze the objective function at those corner points.

objective function is c = 30x + 40y

at (45,75), you get 30*45 + 40*75 = 4350

at (45,60), you get 30*45 + 40*60 = 3750

at (59,32), you get 30*59 + 40*32 = 3050

at (88,32), you get 30*88 + 40*32 = 3920

your minimum cost is when x = 59 and y = 32.

your minimum cost is 3050.

all your constraints need to be satisfied as well.

x + y <= 120 becomes 59 + 32 <= 120 which becomes 91 <= 120 which is true.

x >= 45 becomes 59 >= 45 which is true.

y >= 32 becomes 32 >= 32 which is true.

x and y >= 0 are automatically satisfied when x >= 45 and y >= 32.

20x + 10y >= 1500 becomes 59*20 + 32*10 >= 1500 which becomes 1180 + 320 >= 1500 which becomes 1500 >= 1500 which is true.

all the constraints are satisfied.

your solution is that the minimum cost is 3050.